6. Find the result of multiplying a complex number by its conjugate
Answer #1

2 = atib 121= √ a² +6² Z = a-ib Izl= Ja²+(-b?=Ja²7b2 Z) = (atib) (a-ib) -= az - lab tiab -lib) ² =a2 -(-6?) -= a² +62 & Izz =